OpenStudy (anonymous):

Find the center and the radius of the circle: x^2+y^2-12x+6y=19

OpenStudy (anonymous):

group the similar values together (x^2-12x) + (y^2 +6y) = 19

OpenStudy (anonymous):

take out what you can so x(x-12) + y(y+6) = 19

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

standard form of that circle is (x-6)^2 + (y+3)^2=64